Two vectors, 300 and 400, have a common origin and are 60° apart. Find the magnitude and the direction of the resultant using a.) The cosine law and b.) The i and j unit vectors


1
Expert's answer
2021-03-18T00:19:10-0400

a) Let "a=300", "b=400" and angle between them "\beta=180^{\circ}-60^{\circ}=120^{\circ}". Then, according to the cosine law, we get:


"R^2=a^2+b^2-2abcos\beta,""R=\sqrt{a^2+b^2-2abcos\beta},""R=\sqrt{(300)^2+(400)^2-2\cdot300\cdot400\cdot cos120^{\circ}}=608.3"

We can find the direction as follows:


"b^2=a^2+R^2-2aRcos\alpha,""\alpha=cos^{-1}{\dfrac{a^2+R^2-b^2}{2aR}},""\alpha=cos^{-1}{\dfrac{(300)^2+(608.3)^2-(400)^2}{2\cdot300\cdot608.3}},""\alpha=34.7^{\circ}.""\theta=60^{\circ}-\alpha=60^{\circ}-34.7^{\circ}=25.3^{\circ}."


b) Let's find "x" and "y" components of the resultant:


"R_x=400i+(300\cdot cos60^{\circ})i=550i,""R_y=0j+(300\cdot sin60^{\circ})j=260j."

Then, we can write the resultant in unit vector notation as follows:


"R=550i+260j."

We can find the magnitude of the resultant from the Pythagorean theorem:


"R=\sqrt{R_x^2+R_y^2}=\sqrt{(550)^2+(260)^2}=608.3"

We can find the direction from the geometry:


"\theta=cos^{-1}(\dfrac{R_x}{R})=cos^{-1}(\dfrac{550}{608.3})=25.3^{\circ}."